# Large Language Models based ASR Error Correction for Child Conversations

### v1

@@ 1 Introduction
Large Language Models (LLMs), trained on massive amounts of text data, have gained substantial attention in natural language processing through their advanced capabilities in processing large volumes of input data and generating sophisticated inferences and responses [15]. While initially developed for text-related tasks, these models have shown promising applications in ASR systems. For example, several studies [16, 17, 18] have shown that LLMs can perform ASR tasks when integrated with audio or speech encoders. Ogawa et al. [19] used LLMs to improve ASR transcriptions by rescoring ASR hypotheses using conversational contexts.
LLMs’ capability to process language structure, context, and semantic relationships enables them to effectively correct ASR errors by considering both narrow linguistic patterns and broader semantic context [15]. Recent works [20, 21] have demonstrated the effectiveness of leveraging LLMs for ASR error correction by incorporating N-best hypotheses from ASR systems. This allows the models to select and refine the most probable transcription. However, to our best knowledge, there is limited work that examines LLMs for ASR error correction in conversational settings, particularly in child-inclusive contexts. Additionally, these works have mainly focused on improving zero-shot ASR outputs, while LLMs’ capabilities to improve fine-tuned ASR outputs remain unclear.

### v2

@@ Abstract
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
Large Language Models (LLMs) have gained substantial attention in natural language processing through their advanced capabilities in processing large volumes of input data and generating sophisticated inferences and responses [15]. While initially developed for text-related tasks, these models have shown promising applications in ASR systems. For example, several studies [16, 17, 18] have shown that LLMs can perform ASR tasks when integrated with audio or speech encoders. Ogawa et al. [19] used LLMs to improve ASR transcriptions by rescoring ASR hypotheses using conversational contexts.
LLMs’ capability to process language structure, context, and semantic relationships enables them to effectively correct ASR errors by considering both narrow linguistic patterns and broader semantic context [15]. Recent works [20, 21] have demonstrated the effectiveness of leveraging LLMs for ASR error correction by selecting and refining the most probable transcription using N-best hypotheses from ASR systems. However, to our knowledge, limited work examines LLMs for ASR error correction in conversational settings, particularly in child-inclusive contexts. Additionally, these works have mainly focused on improving zero-shot ASR outputs, while LLMs’ capabilities to improve fine-tuned ASR outputs remain unclear.